After reviewing a movie, 400 people rated the movie as excellent, good, or fair. the data are given. Construct a pie chart. Excellent- 80
Good-200
Fair-120

Answer by stanbon(75887)

After reviewing a movie, 400 people rated the movie as excellent, good, or fair. the data are given. Construct a pie chart.
Convert the categories to percentage:
Excellent- 80-------80/400 = 20%
Good-200- 200------200/400 = 50%
Fair-120- 120------120/400 = 30%
-------
Draw a circle. Convert each percentage into a pie slice
50% is top half
30% and 20% occupy the lower half.
======================================
Cheers,
Stan H.
===============
----